Solve the triangle a=4 b=3 c=6
Answer by stanbon(75887)

Solve the triangle a=4 b=3 c=6
----
a^2 = b^2+c^2-2bc*Cos(A)
Let A be the angle opposite a = 4
------------------------
Then::
Cos(A) = [4^2 - 3^2-6^2]/(-2*3*6) = 0.8056
A = arcCos(0.8056) = 36.33 degrees
------------
Do the same with B.
----
Then C = 180-(A+B)
----
Cheers,
Stan H.
---------------------